# QMUL-PH-25-01

### v1

@@ 3.3 General expression for ${\cal E}^{w}_{i,j}(\tau)$
where the integer $n$ is related to the eigenvalue $s$ via $n=s-1=3i+j$. We then use (3.36) to derive
@@ 4.1 Fourier mode expansion
where $q=e^{2\pi i\tau}$ and $\bar{q}=e^{-2\pi i\bar{\tau}}$. The mode number is given by $k=n-m$ and modes with $k>0$ represent the contributions of instantons while modes with $k<0$ are contributions of anti-instantons. It is useful to separate the terms with $k=0$ into purely perturbative contributions, coming from the $m=n=0$ term, and instanton/anti-instanton contributions where $m=n>0$.
@@ 4.2 $L$-values and generalised Eisenstein series
$=(-1)^{s}\frac{\pi^{s}i}{\Gamma(s)}\tau_{2}^{1-s}\int_{\tau}^{i\infty}(\tau-\tau_{1})^{s-1}(\bar{\tau}-\tau_{1})^{s-1}\,\Delta(\tau)\,{\rm d}\tau_{1}+{\rm c.c.}\,,$ (4.15)
@@
$\dim\mathcal{S}_{2s}=\left\{\begin{array}[]{lc}\left\lfloor\frac{2s}{12}\right\rfloor-1&2s\equiv 2\,{\rm mod}\,12\,,\\[5.69054pt] \left\lfloor\frac{2s}{12}\right\rfloor\phantom{-1}&\mbox{otherwise}\,,\end{array}\right.$ (4.17)
@@
where $w=s_{1}+s_{2}$. From the spectrum (4.13) it is easy to see that since $w=s_{1}+s_{2}\geq s+2$ it follows that the ratio of $L$-values in the coefficients (4.19) always consists of an odd critical $L$-value in the denominator and an odd non-critical $L$-value in the numerator where, for a cusp form $\Delta\in\mathcal{S}_{2s}$, the critical $L$-values are the numbers $\Lambda(\Delta,t)$ with $t\in\{1,\ldots,2s-1\}$. The general expression for the coefficients $\kappa_{\Delta,s,s_{1},s_{2}}$ is not determined at this stage but they can be evaluated on a case by case basis by evaluating the multiple modular values associated with the particular solution part of (4.16) and they are generically in the number field associated with $S_{2s}$. [^11] Whenever $\dim\mathcal{S}_{2s}=1$ this number field actually coincides with the rational numbers. However, when $\dim\mathcal{S}_{2s}>1$ that is no longer true. For example, with $242s=$ 24 we have $\dim\mathcal{S}_{24}=2$ and the Fourier coefficients of the two Hecke eigenforms basis elements of $\mathcal{S}_{24}$ belong to the number field $\mathbb{Q}[\sqrt{144169}]$. by the Fourier coefficients $a(n)$ of $\Delta(\tau)=\sum_{n=1}^{\infty}a(n)q^{n}$. However, we will shortly propose a strongly motivated conjectural expression for the coefficients in (4.19) based on the expression (4.41) for these coefficients in the context of SMFs, which will be discussed in the next sub-section.
@@
with $w=s_{1}+s_{2}$ and where $\widetilde{C}_{s;s_{1},s_{2}}$ is the rational number defined in (3.50). The number field content for the coefficient $\lambda_{\Delta}(s;s_{1},s_{2})$ in the expression (4.21) is entirely carried by the Petersson norm $\langle\Delta,\Delta\rangle$ in the denominator, since $\tilde{\kappa}_{\Delta,s,s_{1},s_{2}}$ in (4.22) is manifestly rational for the spectrum (4.13). Furthermore, for this spectrum $\lambda_{\Delta}(s;s_{1},s_{2})$ always contains the product of an even critical $L$-value and an odd non-critical $L$-value.
@@
We now consider the GES solution to (4.1) with spectrum [^12] As mentioned in footnote 7, the case with either $s_{1}$ or $s_{2}$ equal to $\frac{1}{2}$ is slightly different from the other half-integer cases since the Fourier zero-mode sector does contain a term proportional to $\log(\tau_{2})$. Although our present discussion encompasses this case as well, these particular modular invariant functions do not seem to be relevant when discussing SMFs.
@@ 4.3 Absence of $L$-values in theta-lifted local Maass functions
Using the results of section 3, we see that if we want to obtain linear combinations of GESs with parameters in the spectrum (4.13) relevant for MGFs we have to consider the theta lift (2.32), or equivalently (2.34), for $w\geq 4$ and $21\leq s\leq w-$ 2 with the same parity. As a consequence we see from (3.13) that the integrand (2.35) reduces to a symmetric polynomial in $t_{i}$ of degree $w-3$.
It is then straightforward to perform the integration (2.34) using the basic integral representation of a MGF given in (2.40). We deduce that the theta lift (2.32) of a local Maass function with $w\geq 4$ and $21\leq s\leq w-$ 2 produces a rational linear combination of MGFs ($C_{a,b,c}(\tau)$ with $w=a+b+c$), that solves an inhomogeneous Laplace equation with eigenvalue $s$ and source terms given by bilinears, $E(s_{1};\tau)E(s_{2};\tau)$, in non-holomorphic Eisenstein series with $w=s_{1}+s_{2}$ and possibly a multiple of the single non-holomorphic Eisenstein series $E(w;\tau)$.
@@
Note that this matches the total number of functions $B^{w}_{i,j}$ defined in (2.35) consistent with the spectrum (4.13) given that $s=3i+j+1$. Hence we see that the theta lift off local Maass functions with $w\geq 3$ and $21\leq s\leq w-$ 2 results in the vector space of $C_{a,b,c}(\tau)$ conveniently arranged in terms of eigenvectors of the Laplace operator.
@@
Since $\dim\mathcal{S}_{2s}=0$ for $s<6$ and since $s<w-2$ it follows that at weight $w=8$ and at any weight $w\geq 10$, the number of independent GESs is strictly larger than that of $C_{a,b,c}(\tau)$. More precisely, the number of GESs at weight $w$ and eigenvalue $s$ in the MGF spectrum (4.13) that cannot be written in terms of MGFs equals the number of holomorphic cusp forms at modular weight $2s$ (which is given in (4.17)).
@@
To extract a particular MGF from the generating series we simply need to extract the coefficient of certain non-commutative sign-post variables, which furnish a representation of an important mathematical object called Tsunogai’s derivation algebra [42]. As a consequence of this fact, we find that MGFs are related only to specific linear combinations of equivariant iterated Eisenstein integrals, namely those that follow from the so-called Tsunogai’s relations [42, 43], see also [7, 41, 10, 38, 44]. While the modular completion of iterated integrals of holomorphic Eisenstein series necessitates the addition of iterated integrals of holomorphic cusp forms [5, 7], such as for the depth-two examples (4.18), the linear combinations of equivariant iterated integrals selected by Tsunogai’s relations and hence corresponding to MGFs are expressible in terms of iterated integrals involving solely holomorphic Eisenstein series (and their complex conjugates) and in addition the periods arising in this construction are conjecturally restricted solely to single-valued MZVs.
As a concrete consequence of this structure, we find that the theta lift of modular local polynomials corresponding to the spectrum (4.13) produces precisely ths linear combinations of GESs in (4.47)-(4.48), or equivalently of depth-two equivariant iterated Eisenstein integrals for which the cusp form contributions have dropped out. We expect that a generalisation of (2.32) to a higher-dimensional lattice sum will produce higher-loop MGFs and hence the corresponding linear combinations of higher-depth iterated Eisenstein integrals selected by Tsunogai’s relations.
@@
Since the modular local polynomial $u^{i}v^{j}$ transforms under $U$ with modular weight $-2(3i+j)=-2(s-1)$ and given that $\Delta\in\mathcal{S}_{2s}$ we see that $\Delta(\tau)\,u^{i}v^{j}{\rm d}\tau$ is invariant under $U$. We deduce that under the change of integration variables $\tau=U\cdot\tau^{\prime}$ equation (4.59) becomes
@@
Furthermore, a quick counting argument shows that any rational linear combination of GESs, $\mathcal{E}(s;s_{1},s_{2};\tau)$, at fixed weight $w=s_{1}+s_{2}$ and eigenvalue $s$ with parameters $(s,s_{1},s_{2})$ in either (4.13) or (4.24) and for which the cusp form contribution cancels must be a linear combination of theta lifts ${\cal E}_{i,j}^{w}(\tau)$ with $s=3i+j+1$. For the spectrum of parameters (4.13) it is easy to see that at fixed eigenvalue $s$ we have $\lfloor\frac{s}{2}\rfloor$ linearly independent GESs at fixed weight $w=s_{1}+s_{2}$. Similarly, for the spectrum of parameters (4.24) at fixed eigenvalue $s$ the largest number of GESs is found at weight $w=s-1$ where there are again $\lfloor\frac{s}{2}\rfloor$ linearly independent GESs. In both cases we can write a general rational linear combination of GESs and impose cancellation of the $L$-values, using either the conjectural expression (4.21) or (4.41), and we obtain a system of ${\rm dim}\,\mathcal{S}_{2s}$ equations [^13] Since we are working with rational linear combinations of GESs and the cusp form coefficients (4.21) and (4.41) take values in the number field associated with $\mathcal{S}_{2s}$ we deduce that a single linear equation with number field coefficients becomes a system of ${\rm dim}\,\mathcal{S}_{2s}$ equations with rational coefficients. in $\lfloor\frac{s}{2}\rfloor$ variables. From the dimension formula (4.17) we see that this system of equations has exactly $\lfloor\frac{s-1}{3}\rfloor$ linearly independent solutions, which is precisely the number of ${\cal E}_{i,j}^{w}(\tau)$ with $s=3i+j+1$.
@@
The first value of $i$ for which the identity $\mathcal{K}_{i,0}(\Delta_{6i+2})=0$ becomes non-trivial is $i=3$ for which we must consider the unique Hecke normalised cusp form of weight $206i+2=$ 20 and, after using the $L$-function reflection identity (C.8), equation (4.62) for $i=3$ reduces to
$-\frac{\Lambda(\Delta_{20},10)}{36}+\frac{\Lambda(\Delta_{20},12)}{60}=0\,,$ (4.63)
@@ 5.3 S-dual modular forms
Explicit examples discussed in [46] include the maximal $U(1)$-violating scattering amplitudes of four gravitons and $h$ axion-dilatons excitations in type IIB superstring theory, which are described by modular forms with holomorphic and anti-holomorphic weights $(h,-h)$, here denoted by ${\cal E}_{h}(s;s_{1},h_{1};s_{2},h_{2};\tau)$. In general, such modular forms transform as
@@ Appendix B Modular local polynomials
We also see that $[-1,1,0]\in\mathcal{Q}_{1}$ with set
@@
Since $\tilde{\rho}\in\mathcal{F}_{\Gamma_{0}(2)}$ we know already that $10<x<$ 1 and $x^{2}+y^{2}>x$ so we deduce
@@
To satisfy (B.10) given this constraint we must have $-b>a+c\geq 0$ hence the discriminant must satisfy
@@ Appendix C Holomorphic cusp forms and their $L$-values
$[T_{n}\Delta](\tau)=n^{k-1}\sum_{d|n}d^{-k}\sum_{b=0}^{d-1}\Delta\left(\frac{n\tau+bd}{d^{2}}\right)\,.$ (C.6)
A preferred basis for $\mathcal{S}_{k}$ can be constructed by considering the basis of Hecke eigenforms which simultaneously diagonalise all $T_{n}$. Importantly, the eigenvalues of a Hecke eigenform $\Delta(\tau)=\sum_{n=1}^{\infty}a(n)q^{n}\ \in\mathcal{S}_{k}$ are precisely given by its Fourier $q$-series coefficients, i.e. we have $[T_{n}\Delta](\tau)=a(n)\Delta(\tau)$ so that from (C.6) we see that a Hecke eigenform is automatically Hecke normalised.
@@
which is a degree $k-2$ polynomial in the variable $X$. We can use the explicit $q$-series representation $\Delta(\tau)=\sum_{n=1}^{\infty}a(n)q^{n}$ with $q=\exp(2\pi i\tau)$ to show that
@@
where $B_{n}$ is the $n^{th}$ Bernoulli number and $q$ and $r$ are any even integers such that $q+r=k$ and $24\leq r\leq\frac{k}{2}-$ 2. The function $\kappa(j,q,r)$ takes value in the number field generated by the Fourier coefficients of the Hecke eigenforms of $\mathcal{S}_{k}$ and it is given by
@@
with $\alpha_{k}(n)$ given in (C.3). The $({\rm dim}\,\mathcal{S}_{k})\times({\rm dim}\,\mathcal{S}_{k})$ matrix $H$ is defined by $H_{nj}\coloneqq a^{(j)}(n)$, i.e. the entry in the $n^{th}$ row and $j^{th}$ column of the matrix $H$ is precisely the $n^{th}$ $q$-series coefficient of the $j^{th}$ Hecke eigenform, i.e. $\Delta^{(j)}(\tau)=\sum_{n=1}^{\infty}a^{(j)}(n)q^{n}\in\mathcal{S}_{k}$. A particularly simple expression can be derived when we specialise (C.16) to the cases where ${\rm dim}\,\mathcal{S}_{k}=1$, i.e. $k\in\{12,16,20,22,26\}$. In this case the $1\times 1$ matrix $H$ reduces to $1$ from the Hecke normalisation, and if we furthermore choose $r=4,q=k-4$ we find that (C.16) reduces to
[…]

### v2

@@ 3.3 General expression for ${\cal E}^{w}_{i,j}(\tau)$
where the integer $n$ is related to the eigenvalue $s$ via n=s1=3i+jn=s-1=3i+j. We then use (3.36) to derive
@@ 4.1 Fourier mode expansion
where $q=e^{2\pi i\tau}$ and $\bar{q}=e^{-2\pi i\bar{\tau}}$. The mode number is given by k=nmk=n-m and modes with $k>0$ represent the contributions of instantons while modes with $k<0$ are contributions of anti-instantons. It is useful to separate the terms with $k=0$ into purely perturbative contributions, coming from the $m=n=0$ term, and instanton/anti-instanton contributions where $m=n>0$.
@@ 4.2 $L$-values and generalised Eisenstein series
$=(-1)^{s}\frac{\pi^{s}i}{\Gamma(s)}\tau_{2}^{1-s}\int_{\tau}^{i\infty}(\tau-\tau^{\prime})^{s-1}(\bar{\tau}-\tau^{\prime})^{s-1}\,\Delta(\tau^{\prime})\,{\rm d}\tau^{\prime}+{\rm c.c.}\,$, (4.15)
@@
[…]
@@
where $w=s_{1}+s_{2}$. From the spectrum (4.13) it is easy to see that since $w=s_{1}+s_{2}\geq s+2$ it follows that the ratio of $L$-values in the coefficients (4.19) always consists of an odd critical $L$-value in the denominator and an odd non-critical $L$-value in the numerator where, for a cusp form $\Delta\in\mathcal{S}_{2s}$, the critical $L$-values are the numbers $\Lambda(\Delta,t)$ with $t\in\{1,\ldots,2s-1\}$. The general expression for the coefficients $\kappa_{\Delta,s,s_{1},s_{2}}$ is not determined at this stage but they can be evaluated on a case by case basis by evaluating the multiple modular values associated with the particular solution part of (4.16) and they are generically in the number field associated with $\mathcal{S}_{2s}$. [^12] Whenever $\dim\mathcal{S}_{2s}=1$ this number field actually coincides with the rational numbers. However, when $\dim\mathcal{S}_{2s}>1$ that is no longer true. For example, with $2s=24$ we have $\dim\mathcal{S}_{24}=2$ and the Fourier coefficients of the two Hecke eigenforms basis elements of $\mathcal{S}_{24}$ belong to the number field $\mathbb{Q}[\sqrt{144169}]$. by the Fourier coefficients $a(n)$ of $\Delta(\tau)=\sum_{n=1}^{\infty}a(n)q^{n}$. However, we will shortly propose a strongly motivated conjectural expression for the coefficients in (4.19) based on the expression (4.41) for these coefficients in the context of SMFs, which will be discussed in the next sub-section.
@@
with $w=s_{1}+s_{2}$ and where $\widetilde{C}_{s;s_{1},s_{2}}$ is the rational number defined in (3.50). Our choice of normalisation ensures that the overall coefficient $\tilde{\kappa}_{\Delta,s,s_{1},s_{2}}$ in (4.22) is manifestly rational for the spectrum (4.13). Furthermore, for this spectrum $\lambda_{\Delta}(s;s_{1},s_{2})$ always contains the product of an even critical $L$-value and an odd non-critical $L$-value. Lastly, while the particular results provided in [9, 10] have the completed $L$-values multiplied by an overall coefficient which explicitly belongs to the number field associated with $\mathcal{S}_{2s}$, we see that in our expression (4.21) the dependence on the number field is implicit and it is entirely captured by the Petersson norm $\langle\Delta,\Delta\rangle$ in the denominator, as we see in Rankin’s expression (C.16) for the Petersson norm.
@@
We now consider the GES solution to (4.1) with spectrum [^13] As mentioned in footnote 8, the case with either $s_{1}$ or $s_{2}$ equal to $\frac{1}{2}$ is slightly different from the other half-integer cases since the Fourier zero-mode sector does contain a term proportional to $\log(\tau_{2})$. Although our present discussion encompasses this case as well, these particular modular invariant functions do not seem to be relevant when discussing SMFs.
@@ 4.3 Absence of $L$-values in theta-lifted local Maass functions
Using the results of section 3, we see that if we want to obtain linear combinations of GESs with parameters in the spectrum (4.13) relevant for MGFs we have to consider the theta lift (2.32), or equivalently (2.34), for $w\geq 4$ and $1\leq s\leq w-2$ with the same parity. As a consequence we see from (3.13) that the integrand (2.35) reduces to a symmetric polynomial in $t_{i}$ of degree w3w-3.
It is then straightforward to perform the integration (2.34) using the basic integral representation of a MGF given in (2.40). We deduce that the theta lift (2.32) of a local Maass function with $w\geq 4$ and $1\leq s\leq w-2$ produces a rational linear combination of MGFs ($C_{a,b,c}(\tau)$ with $w=a+b+c$), that solves an inhomogeneous Laplace equation with eigenvalue $s$ and source terms given by bilinears, $E(s_{1};\tau)E(s_{2};\tau)$, in non-holomorphic Eisenstein series with $w=s_{1}+s_{2}$ and possibly a multiple of the single non-holomorphic Eisenstein series $E(w;\tau)$.
@@
Note that this matches the total number of functions $B^{w}_{i,j}$ defined in (2.35) consistent with the spectrum (4.13) given that $s=3i+j+1$. Hence we see that the theta lift off local Maass functions with $w\geq 3$ and $1\leq s\leq w-2$ results in the vector space of $C_{a,b,c}(\tau)$ conveniently arranged in terms of eigenvectors of the Laplace operator.
@@
Since $\dim\mathcal{S}_{2s}=0$ for $s<6$ and since s<w2s<w-2 it follows that at weight $w=8$ and at any weight $w\geq 10$, the number of independent GESs is strictly larger than that of $C_{a,b,c}(\tau)$. More precisely, the number of GESs at weight $w$ and eigenvalue $s$ in the MGF spectrum (4.13) that cannot be written in terms of MGFs equals the number of holomorphic cusp forms at modular weight $2s$ (which is given in (4.17)).
@@
To extract a particular MGF from the generating series we simply need to extract the coefficient of certain non-commutative sign-post variables, which conjecturally furnish a representation of an important mathematical object called Tsunogai’s derivation algebra [42]. As a consequence of this fact, we find that MGFs are conjecturally related only to specific linear combinations of equivariant iterated Eisenstein integrals, namely those that follow from the so-called Tsunogai’s relations [42, 43], see also [7, 41, 10, 38, 44]. While the modular completion of iterated integrals of holomorphic Eisenstein series necessitates the addition of iterated integrals of holomorphic cusp forms [5, 7], such as for the depth-two examples (4.18), the linear combinations of equivariant iterated integrals selected by Tsunogai’s relations and hence conjecturally corresponding to MGFs are expressible in terms of iterated integrals involving solely holomorphic Eisenstein series (and their complex conjugates) and in addition the periods arising in this construction are conjecturally restricted solely to single-valued MZVs.
As a concrete consequence of this structure, we find that the theta lift of modular local polynomials corresponding to the spectrum (4.13) produces precisely the linear combinations of GESs in (4.47)-(4.48), or equivalently of depth-two equivariant iterated Eisenstein integrals for which the cusp form contributions have dropped out. We expect that a generalisation of (2.32) to a higher-dimensional lattice sum will produce higher-loop MGFs and hence the corresponding linear combinations of higher-depth iterated Eisenstein integrals selected by Tsunogai’s relations.
@@
Since the modular local polynomial $u^{i}v^{j}$ transforms under $U$ with modular weight 2(3i+j)=2(s1)-2(3i+j)=-2(s-1) and given that $\Delta\in\mathcal{S}_{2s}$ we see that $\Delta(\tau)\,u^{i}v^{j}{\rm d}\tau$ is invariant under $U$. We deduce that under the change of integration variables $\tau=U\cdot\tau^{\prime}$ equation (4.59) becomes
@@
Furthermore, a quick counting argument shows that any rational linear combination of GESs, $\mathcal{E}(s;s_{1},s_{2};\tau)$, at fixed weight $w=s_{1}+s_{2}$ and eigenvalue $s$ with parameters $(s,s_{1},s_{2})$ in either (4.13) or (4.24) and for which the cusp form contribution cancels must be a linear combination of theta lifts ${\cal E}_{i,j}^{w}(\tau)$ with $s=3i+j+1$. For the spectrum of parameters (4.13) it is easy to see that at fixed eigenvalue $s$ we have $\lfloor\frac{s}{2}\rfloor$ linearly independent GESs at fixed weight $w=s_{1}+s_{2}$. Similarly, for the spectrum of parameters (4.24) at fixed eigenvalue $s$ the largest number of GESs is found at weight w=s1w=s-1 where there are again $\lfloor\frac{s}{2}\rfloor$ linearly independent GESs. In both cases we can write a general rational linear combination of GESs and impose cancellation of the $L$-values, using either the conjectural expression (4.21) or (4.41), and we obtain a system of ${\rm dim}\,\mathcal{S}_{2s}$ equations [^14] Since we are working with rational linear combinations of GESs and the cusp form coefficients (4.21) and (4.41) take values in the number field associated with $\mathcal{S}_{2s}$ we deduce that a single linear equation with number field coefficients becomes a system of ${\rm dim}\,\mathcal{S}_{2s}$ equations with rational coefficients. in $\lfloor\frac{s}{2}\rfloor$ variables. From the dimension formula (4.17) we see that this system of equations has exactly $\lfloor\frac{s-1}{3}\rfloor$ linearly independent solutions, which is precisely the number of ${\cal E}_{i,j}^{w}(\tau)$ with $s=3i+j+1$.
@@
The first value of $i$ for which the identity $\mathcal{K}_{i,0}(\Delta_{6i+2})=0$ becomes non-trivial is $i=3$ for which we must consider the unique Hecke normalised cusp form of weight $6i+2=20$ and, after using the $L$-function reflection identity (C.8), equation (4.62) for $i=3$ reduces to
$\frac{\Lambda(\Delta_{20},10)}{36}+\frac{\Lambda(\Delta_{20},12)}{60}=0\,$, (4.63)
@@ 5.3 S-dual modular forms
Explicit examples discussed in [46] include the maximal $U(1)$-violating scattering amplitudes of four gravitons and $h$ axion-dilatons excitations in type IIB superstring theory, which are described by modular forms with holomorphic and anti-holomorphic weights (h,h)(h,-h), here denoted by ${\cal E}_{h}(s;s_{1},h_{1};s_{2},h_{2};\tau)$. In general, such modular forms transform as
@@ Appendix B Modular local polynomials
We also see that $1,1,0]\in\mathcal{Q}_{1}$ with set
@@
Since $\tilde{\rho}\in\mathcal{F}_{\Gamma_{0}(2)}$ we know already that $0<x<1$ and $x^{2}+y^{2}>x$ so we deduce
@@
To satisfy (B.10) given this constraint we must have $b>a+c\geq 0$ hence the discriminant must satisfy
@@ Appendix C Holomorphic cusp forms and their $L$-values
$T_{n}\Delta](\tau)=n^{k-1}\sum_{d|n}d^{-k}\sum_{b=0}^{d-1}\Delta\left(\frac{n\tau+bd}{d^{2}}\right)\,.$ (C.6)
A preferred basis for $\mathcal{S}_{k}$ can be constructed by considering the basis of Hecke eigenforms which simultaneously diagonalise all $T_{n}$. Importantly, the eigenvalues of a Hecke eigenform $\Delta(\tau)=\sum_{n=1}^{\infty}a(n)q^{n}\ \in\mathcal{S}_{k}$ are precisely given by its Fourier $q$-series coefficients, i.e. we have $T_{n}\Delta](\tau)=a(n)\Delta(\tau)$ so that from (C.6) we see that a Hecke eigenform is automatically Hecke normalised.
@@
which is a degree k2k-2 polynomial in the variable $X$. We can use the explicit $q$-series representation $\Delta(\tau)=\sum_{n=1}^{\infty}a(n)q^{n}$ with $q=\exp(2\pi i\tau)$ to show that
@@
where $B_{n}$ is the $n^{th}$ Bernoulli number and $q$ and $r$ are any even integers such that $q+r=k$ and $4\leq r\leq\frac{k}{2}-2$. The function $\kappa(j,q,r)$ takes value in the number field generated by the Fourier coefficients of the Hecke eigenforms of $\mathcal{S}_{k}$ and it is given by
@@
with $\alpha_{k}(n)$ given in (C.3). The $({\rm dim}\,\mathcal{S}_{k})\times({\rm dim}\,\mathcal{S}_{k})$ matrix $H$ is defined by $H_{nj}\coloneqq a^{(j)}(n)$, i.e. the entry in the $n^{th}$ row and $j^{th}$ column of the matrix $H$ is precisely the $n^{th}$ $q$-series coefficient of the $j^{th}$ Hecke eigenform, i.e. $\Delta^{(j)}(\tau)=\sum_{n=1}^{\infty}a^{(j)}(n)q^{n}\in\mathcal{S}_{k}$. A particularly simple expression can be derived when we specialise (C.16) to the cases where ${\rm dim}\,\mathcal{S}_{k}=1$, i.e. $k\in\{12,16,20,22,26\}$. In this case the $1\times 1$ matrix $H$ reduces to $1$ from the Hecke normalisation, and if we furthermore choose r=4,q=k4r=4,q=k-4 we find that (C.16) reduces to
${\langle\Delta,\Delta\rangle}=-\frac{15\ 2^{5-k}(k-4)B_{k}\,\Lambda(\Delta,k-4)\Lambda(\Delta,k-1)}{\big{[}(k-4)B_{k}-(k+120B_{k})B_{k-4}\big{]}}\,.$ (C.18)